OpenStudy (aravindg):

is there any easy way to findcross product AxAxA rather than finding product AxA and multi[lying by A??

OpenStudy (2bornot2b):

I would like the answer to this question

OpenStudy (radar):

Since A has implied an exponent of 1. Simply ad the exponents. AxAxA=A^3 \[a ^{2}\times a ^{1}=a ^{3}\] etc.

OpenStudy (2bornot2b):

I think asker is talking about vectors, not scalars.

OpenStudy (aravindg):

lol not that its cross product

OpenStudy (aravindg):

no

OpenStudy (aravindg):

i am asking cross product in functions

OpenStudy (radar):

Never mind lol.

OpenStudy (jamesj):

AxA = 0 because the A is parallel to itself. Hence AxAxA = 0 also

OpenStudy (jamesj):

Make sense?

OpenStudy (aravindg):

no i didnt ask that

OpenStudy (jamesj):

What's your question then.

OpenStudy (aravindg):

i am asking abot ordered pairs

OpenStudy (jamesj):

You mean this identity? \[ A \times B \times C = B(A\cdot C) - C(A \cdot B) \]

OpenStudy (jamesj):

\[ A \times (B \times C) \]